# Coulomb Branch Operator Algebras and Universal Selection Rules for $\mathcal{N}=2$ SCFTs

### v1

@@ Abstract
Coulomb branches of vacua are the most universal moduli spaces that arise in local unitary interacting 4d $\mathcal{N}=2$ superconformal field theories (SCFTs). In these theories, $1/2$-BPS primaries parameterize the Coulomb branches and form (anti-)chiral rings. We define the notion of a Coulomb branch operator algebra, $\mathcal{A}_{\mathcal{C}}$, that contains these chiral and anti-chiral rings along with infinitely many more operators and products that are less protected by supersymmetry. Using a universal symmetry, $\mathcal{I}\cong\mathbb{Z}_{2}$, that arises from studying the superconformal group, we give $\mathcal{I}$ selection rules for $\mathcal{A}_{\mathcal{C}}$ and, more generally, for arbitrary products in the local operator algebra of any 4d $\mathcal{N}=2$ SCFT. Defining the notion of a “Coulombic” SCFT, we propose explanations for certain phenomena in a 4d/2d correspondence involving 4d $\mathcal{N}=2$ theories and 2d vertex operator algebras (VOAs). Finally, by considering a deformation of $\mathcal{I}$, we explore the case of $\mathcal{N}=4$ SCFTs.
@@ 1 Introduction
$\mathcal{A}_{\mathcal{C}}:=\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{N}}(y_{N,a_{N}})\Big{\}}~{},$ (1.2)
where $\mathcal{H}_{L}$ is the space of local operators, [^2] Here and throughout this paper we only discuss genuine local operators (e.g., we do not discuss operators that live at the end of Wilson lines). and $\Upsilon_{b}$ is a multiplet of the 4d $\mathcal{N}=2$ superconformal algebra that sits in some (higher) OPE of $\mathcal{E}$ and / or $\bar{\mathcal{E}}$ multiplets. The above product involves all operators in the corresponding generating $\bar{\mathcal{E}}$ and $\mathcal{E}$ multiplets (and derivatives of these operators). For finite central charge, $N$ is finite (moreover, in all known theories, $N$ corresponds to the “rank” or complex dimension of $\mathcal{M}_{C}$). In writing (1.2), we have defined the following (higher) OPE (note that this OPE is not the chiral ring product)
@@
Naively, this discussion suggests that ^$\hat{\mathcal{B}}_{R}\not\in\mathcal{A}_{\mathcal{C}}$. But this argument also has glaring gaps. For example, we know that $\mathcal{A}_{\mathcal{C}}$ is always strictly larger than $\mathcal{R}_{\mathcal{C}}\cup\overline{\mathcal{R}}_{\mathcal{C}}$ (see footnote 4), and we also know there are theories that have Higgs branches but no continuous flavor symmetry (e.g., see [14]). [^6] More generally, it is conceptually possible for there to exist continuous flavor-neutral ^$\hat{\mathcal{B}}_{R}$ multiplets that are not related to moduli spaces. In fact, we will see later that superconformal selection rules on their own do not forbid certain ^$\hat{\mathcal{B}}_{R}$ multiplets from appearing in $\mathcal{A}_{\mathcal{C}}$.
@@
To set up these results, we begin by noting the trivial fact that any multiplet of local operators in a 4d $\mathcal{N}=2$ SCFT lives in one and only one of the following two sectors
$\mathcal{A}_{\mathcal{O},\pm}:=\left\{\Upsilon_{a}\in\mathcal{H}_{L}:\mathcal{I}(\Upsilon_{a})=\pm 1\right\}~{},$ (1.8)
where “$\mathcal{I}(\cdots)$” denotes the charge of the enclosed multiplet. Moreover, $\mathcal{A}_{\mathcal{O},+}$ forms a closed operator subalgebra and must be non-trivial in any local 4d $\mathcal{N}=2$ SCFT since
@@
  • The $\hat{\mathcal{C}}_{R(j,j)}$ multiplets have $R,j,\bar{j}\geq 0$ (locality is conjectured to imply $R\geq|j-\bar{j}|-1$ [24]), $r=j-\bar{j}$, and $\Delta=2+2R+j+\bar{j}$. A prominent example of such a multiplet includes the energy-momentum tensor multiplet, $\hat{\mathcal{C}}_{0(0,0)}$, which is present in any local 4d $\mathcal{N}=2$ SCFT (in a Lagrangian theory, the primary takes the schematic form $J\sim{\rm Tr}(\Phi^{\dagger}\Phi)+\sum_{R}q_{R}^{\dagger}q_{R}$). These multiplets satisfy semi-shortening conditions with respect to the chiral and anti-chiral supercharges. In particular, such multiplets are non-chiral and must exist in any operator algebra that contains multiplets and their CPT conjugates, since $T_{\mu\nu}$ must appear in the corresponding OPEs. The previous sentence explains the discussion around (1.5) in footnote 4.
@@
$\mathcal{A}_{\mathcal{C}}:=\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{\bar{N}}}(y_{\bar{N},a_{\bar{N}}})\Big{\}}~{}.$ (2.9)
@@
Given the above definition, it is easy to see that $\mathcal{H}_{L}$ splits into the following two sectors under the action of $\mathcal{I}$ (where the “$\pm$” superscript refers to $\mathcal{I}$ charge)
[…]
$\ \ \mathcal{A}^{\Delta}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in\mathbb{Z}}\Big{\}}~{},$
[…]
$\ \ \bar{\mathcal{C}}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \mathcal{A}^{\Delta}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}\Big{\}}~{}.$
@@ 2.1 Selection rules, the Schur sector, and the associated VOA
[…]
@@ 2.2 What can appear in the Coulomb branch operator algebra?
$\mathcal{A}_{\mathcal{C}}:=\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{\bar{N}}}(y_{\bar{N},a_{\bar{N}}})\Big{\}}~{}.$ (2.24)
@@ 3 Toward an explanation of VOA spins from the Coulomb branch algebra
We can connect our observation to the discussion in the previous sections by noting that, for SCFTs that lack a Higgs branch (in the sense of footnote 19), it is reasonable to conjecture that the Coulomb branch multiplets generate the local operator algebra, $\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$. The heuristic motivation is that if the only moduli space of vacua is the Coulomb branch, then we expect to be able to reconstruct the SCFT at the origin by studying the Coulomb branch effective theory and certain simple deformations thereof. We can think of the statement that $\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$ as a version of this intuition at the origin of the moduli space.
Conjecture 7: All of the theories in the above list are Coulombic. [^22] As discussed in the previous section, the free vector theory is Coulombic. Moreover, we described non-trivial evidence from the $\bar{\mathcal{C}}$ spectrum suggesting that the MAD theory is also Coulombic (see [26, 27] for a more detailed discussion of this evidence and the associated conjecture). In light of theorem 5, the fact that the associated VOA has integer 2d spin is further evidence for the conjecture that the MAD theory is Coulombic. In fact, using the results of [46], we can strengthen this argument. In this latter paper, we showed that the Lee-Yang vacuum character can be understood as arising from the fact that the following MAD normal-ordered product vanishes: $J^{11}_{+\dot{+}}J^{11}_{+\dot{+}}=0$ (where $J^{11}_{+\dot{+}}$ is the highest $\mathfrak{su}(2)_{R}\oplus\mathfrak{so}(4)$-weight component of the $\mathfrak{su}(2)_{R}$ current). Indeed, this result strongly suggests that $\mathcal{A}_{\rm Schur}^{\rm MAD}<\mathcal{A}^{\rm MAD}_{\mathcal{C}}~{}.$ (3.1) More generally, local unitary 4d $\mathcal{N}=2$ SCFTs lacking a Higgs branch (in the sense of footnote 19) are Coulombic (for theories that have a marginal coupling, we assume we work at a generic point on the conformal manifold away from cusps or other special points [^23] Note that at the $g=0$ point of the $(3,2)$ conformal manifold, the theory cannot be Coulombic: the corresponding ^$\hat{\mathcal{B}}_{R}$ multiplets cannot appear in the OPEs of the vector multiplets we add to gauge the $SU(2)$. Moreover, operators like the ^$\hat{\mathcal{B}}_{3}$ multiplet corresponding to the $g=0$ Schur operator $\mathcal{O}_{abc}:=\epsilon_{IJK}\mu_{1}^{I}\mu_{2}^{J}\mu_{3}^{K}$ cannot appear in any of the $\mathcal{A}_{\mathcal{C}}$ algebras of the individual $(A_{1},A_{3})$ SCFTs (here the subscript of $\mu^{I}_{i}$ denotes the particular $(A_{1},A_{3})$ factor out of the three present, and $I,J,K$ are $\mathfrak{su}(2)$ adjoint gauge indices). Indeed, the reason is that each such operator is charged under the individual $\mathfrak{su}(2)$ symmetries of a given $(A_{1},A_{3})$ factor, but we have seen that Coulomb branch operators are uncharged under flavor Lie algebras.).
@@ 4 Theories with $\mathcal{N}>2$ SUSY
$\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=4}$ ::=:= $\Big{\{}L\bar{B}_{1}[2j;0]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{A}_{2}[2j;0]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},$ (4.8)
$\ L\bar{A}_{1}[2j;2\bar{j}\geq 1]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{L}[j;\bar{j}]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}\Big{\}}~{},$
@@
Indeed, otherwise the SCFT in question would be very unusual: multiplets in $\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=4}$ would not contribute to any correlation functions involving $\mathcal{N}=4$ multiplets containing $\mathcal{N}=2$ Coulomb branch and Schur sector multiplets. However, see footnote 26 for potential sources of counterexamples (in some sense, it would be more interesting for this conjecture to be wrong).
@@
$\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=3}$ $=$ $\Big{\{}L\bar{B}_{1}[2j;0]^{(R_{1},R_{2};r_{3}>r_{*}+6)}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{A}_{2}[2j;0]^{(R_{1},R_{2};r_{3}>r_{*})}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},$ (4.14)
@@
$\ L\bar{L}[j;\bar{j}]^{(R_{1},R_{2};r_{3})}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}\Big{\}}~{},$

### v2

@@ Abstract
Coulomb branches of vacua are the most universal moduli spaces that arise in local unitary interacting 4d $\mathcal{N}=2$ superconformal field theories (SCFTs). In these theories, $1/2$-BPS primaries parameterize the Coulomb branches and form (anti-)chiral rings. We define the notion of a Coulomb branch operator algebra, $\mathcal{A}_{\mathcal{C}}$, that contains these chiral and anti-chiral rings along with infinitely many more operators and products that are less protected by supersymmetry. Using a universal symmetry, $\mathcal{I}\cong\mathbb{Z}_{2}$, that arises from studying the superconformal group, we give $\mathcal{I}$ selection rules for $\mathcal{A}_{\mathcal{C}}$ and, more generally, for arbitrary products in the local operator algebra of any 4d $\mathcal{N}=2$ SCFT. Defining the notion of a “Coulombic” SCFT, we propose explanations for certain phenomena in a 4d/2d correspondence involving 4d $\mathcal{N}=2$ theories and 2d vertex operator algebras. Finally, by considering deformations of $\mathcal{I}$, we explore the case of $\mathcal{N}>2$ SCFTs.
@@ 1 Introduction
$\mathcal{A}_{\mathcal{C}}:={\rm Span}\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{N}}(y_{N,a_{N}})\Big{\}}~{},$ (1.2)
where $\mathcal{H}_{L}$ is the space of local operators, [^2] Here and throughout this paper we only discuss genuine local operators (e.g., we do not discuss operators that live at the end of Wilson lines). and $\Upsilon_{b}$ is a multiplet of the 4d $\mathcal{N}=2$ superconformal algebra that sits in some (higher) OPE of $\mathcal{E}$ and / or $\bar{\mathcal{E}}$ multiplets. [^3] Note that, in taking the span in (1.2), we allow for linear combinations involving different component operators appearing in the $\Upsilon_{b}$ multiplets. The above product involves all operators in the corresponding generating $\bar{\mathcal{E}}$ and $\mathcal{E}$ multiplets (and derivatives of these operators). For finite central charge, $N$ is finite (moreover, in all known theories, $N$ corresponds to the “rank” or complex dimension of $\mathcal{M}_{C}$). In writing (1.2), we have defined the following (higher) OPE (note that this OPE is not the chiral ring product)
@@
Naively, this discussion suggests that ^$\hat{\mathcal{B}}_{R}\not\in\mathcal{A}_{\mathcal{C}}$. But this argument also has glaring gaps. For example, we know that $\mathcal{A}_{\mathcal{C}}$ is always strictly larger than $\mathcal{R}_{\mathcal{C}}\cup\overline{\mathcal{R}}_{\mathcal{C}}$ (see footnote 5), and we also know there are theories that have Higgs branches but no continuous flavor symmetry (e.g., see [14]). [^7] More generally, it is conceptually possible for there to exist continuous flavor-neutral ^$\hat{\mathcal{B}}_{R}$ multiplets that are not related to moduli spaces. In fact, we will see later that superconformal selection rules on their own do not forbid certain ^$\hat{\mathcal{B}}_{R}$ multiplets from appearing in $\mathcal{A}_{\mathcal{C}}$.
@@
To set up these results, we begin by noting the trivial fact that any multiplet of local operators in a 4d $\mathcal{N}=2$ SCFT with definite superconformal quantum numbers, $\Upsilon_{a}$, lives in one and only one of the following two sectors
$\mathcal{A}_{\mathcal{O},\pm}:={\rm Span}\left\{\Upsilon_{a}\in\mathcal{H}_{L}:\mathcal{I}(\Upsilon_{a})=\pm 1\right\}~{},$ (1.8)
where “$\mathcal{I}(\cdots)$” denotes the charge of the enclosed multiplet (comments similar to those in footnote 3 apply for linear combinations of operators in $\mathcal{A}_{\mathcal{O},\pm}$). Moreover, $\mathcal{A}_{\mathcal{O},+}$ forms a closed operator subalgebra and must be non-trivial in any local 4d $\mathcal{N}=2$ SCFT since
@@
  • The $\hat{\mathcal{C}}_{R(j,j)}$ multiplets have $R,j,\bar{j}\geq 0$ (locality is conjectured to imply $R\geq|j-\bar{j}|-1$ [24]), $r=j-\bar{j}$, and $\Delta=2+2R+j+\bar{j}$. A prominent example of such a multiplet includes the energy-momentum tensor multiplet, $\hat{\mathcal{C}}_{0(0,0)}$, which is present in any local 4d $\mathcal{N}=2$ SCFT (in a Lagrangian theory, the primary takes the schematic form $J\sim{\rm Tr}(\Phi^{\dagger}\Phi)+\sum_{R}q_{R}^{\dagger}q_{R}$). These multiplets satisfy semi-shortening conditions with respect to the chiral and anti-chiral supercharges. In particular, such multiplets are non-chiral and must exist in any operator algebra that contains multiplets and their CPT conjugates, since $T_{\mu\nu}$ must appear in the corresponding OPEs. The previous sentence explains the discussion around (1.5) in footnote 5.
@@
$\mathcal{A}_{\mathcal{C}}:={\rm Span}\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{\bar{N}}}(y_{\bar{N},a_{\bar{N}}})\Big{\}}~{}.$ (2.9)
@@
Given the above definition, it is easy to see that the superconformal representations split in the following way under the action of $\mathcal{I}$ (as in (1.8), the “$\pm$” superscript refers to the $\mathcal{I}$ charge; note that we also drop CPT-conjugate multiplets below)
$\mathcal{A}_{\mathcal{O},+}$ $\ni$ ¯$\bar{\mathcal{E}}_{r}~{},\ \hat{\mathcal{B}}_{R}|_{R\in\mathbb{Z}}~{},\ \bar{\mathcal{D}}_{R(j,0)}|_{R+j\in\mathbb{Z}}~{},\ \bar{\mathcal{B}}_{R,r(j,0)}|_{R+j\in\mathbb{Z}}~{},\ \hat{\mathcal{C}}_{R(j,\bar{j})}|_{R+j+\bar{j}\in\mathbb{Z}}~{},\ \bar{\mathcal{C}}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in\mathbb{Z}}$ (2.13)
$\ \ \mathcal{A}^{\Delta}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in\mathbb{Z}}~{},$
$\mathcal{A}_{\mathcal{O},-}$ $\ni$ ^$\hat{\mathcal{B}}_{R}|_{R\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \bar{\mathcal{D}}_{R(j,0)}|_{R+j\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \bar{\mathcal{B}}_{R,r(j,0)}|_{R+j\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \hat{\mathcal{C}}_{R(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},$ (2.15)
$\ \ \bar{\mathcal{C}}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \mathcal{A}^{\Delta}_{R,r(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}~{}.$
@@ 2.1 Selection rules, the Schur sector, and the associated VOA
$\mathcal{A}_{\mathcal{O},+}>\mathcal{A}_{\mathcal{O},+}^{\rm Schur}$ $\ni$ ^$\hat{\mathcal{B}}_{R}|_{R\in\mathbb{Z}}~{},\ \bar{\mathcal{D}}_{R(j,0)}|_{R+j\in\mathbb{Z}}~{},\ \hat{\mathcal{C}}_{R(j,\bar{j})}|_{R+j+\bar{j}\in\mathbb{Z}}~{},$ (2.18)
$\mathcal{A}_{\mathcal{O},-}>\mathcal{A}_{\mathcal{O},-}^{\rm Schur}$ $\ni$ ^$\hat{\mathcal{B}}_{R}|_{R\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \bar{\mathcal{D}}_{R(j,0)}|_{R+j\in{1\over 2}\mathbb{Z}_{\rm odd}}~{},\ \hat{\mathcal{C}}_{R(j,\bar{j})}|_{R+j+\bar{j}\in{1\over 2}\mathbb{Z}_{\rm odd}}~{}.$ (2.19)
@@ 2.2 What can appear in the Coulomb branch operator algebra?
$\mathcal{A}_{\mathcal{C}}:={\rm Span}\Big{\{}\Upsilon_{b}\in\mathcal{H}_{L}:\Upsilon_{b}\in\bar{\mathcal{E}}_{r_{1}}^{\times n_{1}}(x_{1,a_{1}})\times\cdots\times\bar{\mathcal{E}}_{r_{N}}^{\times n_{N}}(x_{N,a_{N}})\times\mathcal{E}_{-r_{1}}^{\times\bar{n}_{1}}(y_{1,a_{1}})\times\cdots\times\mathcal{E}_{-r_{N}}^{\times\bar{n}_{\bar{N}}}(y_{\bar{N},a_{\bar{N}}})\Big{\}}~{}.$ (2.24)
@@ 3 Toward an explanation of VOA spins from the Coulomb branch algebra
We can connect our observation to the discussion in the previous sections by noting that, for SCFTs that lack a Higgs branch (in the sense of footnote 20), it is reasonable to conjecture that the Coulomb branch multiplets generate the local operator algebra, $\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$. The heuristic motivation is that if the only moduli space of vacua is the Coulomb branch, then we expect to be able to reconstruct the SCFT at the origin by studying the Coulomb branch effective theory and certain simple deformations thereof. We can think of the statement that $\mathcal{A}_{\mathcal{C}}=\mathcal{A}_{\mathcal{O}}$ as a version of this intuition at the origin of the moduli space.
Conjecture 7: All of the theories in the above list are Coulombic. [^23] As discussed in the previous section, the free vector theory is Coulombic. Moreover, we described non-trivial evidence from the $\bar{\mathcal{C}}$ spectrum suggesting that the MAD theory is also Coulombic (see [26, 27] for a more detailed discussion of this evidence and the associated conjecture). In light of theorem 5, the fact that the associated VOA has integer 2d spin is further evidence for the conjecture that the MAD theory is Coulombic. In fact, using the results of [46], we can strengthen this argument. In this latter paper, we showed that the Lee-Yang vacuum character can be understood as arising from the fact that the following MAD normal-ordered product vanishes: $J^{11}_{+\dot{+}}J^{11}_{+\dot{+}}=0$ (where $J^{11}_{+\dot{+}}$ is the highest $\mathfrak{su}(2)_{R}\oplus\mathfrak{so}(4)$-weight component of the $\mathfrak{su}(2)_{R}$ current). Indeed, this result strongly suggests that $\mathcal{A}_{\rm Schur}^{\rm MAD}<\mathcal{A}^{\rm MAD}_{\mathcal{C}}~{}.$ (3.1) More generally, local unitary 4d $\mathcal{N}=2$ SCFTs lacking a Higgs branch (in the sense of footnote 20) are Coulombic (for theories that have a marginal coupling, we assume we work at a generic point on the conformal manifold away from cusps or other special points [^24] Note that at the $g=0$ point of the $(3,2)$ conformal manifold, the theory cannot be Coulombic: the corresponding ^$\hat{\mathcal{B}}_{R}$ multiplets cannot appear in the OPEs of the vector multiplets we add to gauge the $SU(2)$. Moreover, operators like the ^$\hat{\mathcal{B}}_{3}$ multiplet corresponding to the $g=0$ Schur operator $\mathcal{O}_{abc}:=\epsilon_{IJK}\mu_{1}^{I}\mu_{2}^{J}\mu_{3}^{K}$ cannot appear in any of the $\mathcal{A}_{\mathcal{C}}$ algebras of the individual $(A_{1},A_{3})$ SCFTs (here the subscript of $\mu^{I}_{i}$ denotes the particular $(A_{1},A_{3})$ factor out of the three present, and $I,J,K$ are $\mathfrak{su}(2)$ adjoint gauge indices). Indeed, the reason is that each such operator is charged under the individual $\mathfrak{su}(2)$ symmetries of a given $(A_{1},A_{3})$ factor, but we have seen that Coulomb branch operators are uncharged under flavor Lie algebras.).
@@ 4 Theories with $\mathcal{N}>2$ SUSY
$\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=4}$ $\ni$ $L\bar{B}_{1}[2j;0]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{A}_{2}[2j;0]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},$ (4.8)
$\ L\bar{A}_{1}[2j;2\bar{j}\geq 1]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{L}[j;\bar{j}]^{(R_{1},R_{2},R_{3})}|_{R_{1}+R_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}~{},$
@@
Indeed, otherwise the SCFT in question would be very unusual: multiplets in $\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=4}$ would not contribute to any correlation functions involving $\mathcal{N}=4$ multiplets containing $\mathcal{N}=2$ Coulomb branch and Schur sector multiplets. However, see footnote 27 for potential sources of counterexamples (in some sense, it would be more interesting for this conjecture to be wrong).
@@
$\mathcal{A}_{\mathcal{O},-}^{\mathcal{N}=3}$ $\ni$ $L\bar{B}_{1}[2j;0]^{(R_{1},R_{2};r_{3}>r_{*}+6)}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},\ L\bar{A}_{2}[2j;0]^{(R_{1},R_{2};r_{3}>r_{*})}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j\in\mathbb{Z}_{\rm odd}}~{},$ (4.14)
@@
$\ L\bar{L}[j;\bar{j}]^{(R_{1},R_{2};r_{3})}|_{{4\over 3}R_{1}+{2\over 3}R_{2}+{1\over 3}r_{3}+2j+2\bar{j}\in\mathbb{Z}_{\rm odd}}~{},$